# Iso-length-spectral Hyperbolic Surface Amalgams

### v1

@@ 3.1 Homeomorphic, non-isometric surface amalgams
Figure 7: An illustration for Case 3, with the same closed geodesic as before except with modified dark green segment. Again, $N=3$ and $c_{2}=2$, so $C_{1}(\gamma)=2(2)=C_{2}(\gamma)$. We can check that there are indeed four identical, non-backtracking closed geodesics in each of $X_{1}$ and $X_{2}$.
We now construct an identical transplanted copy of each $\beta_{i}^{j}$, ${\beta^{\prime}}_{i}^{j}=\bigcup\limits_{k=0}^{L}{\gamma^{\prime}}^{j}_{i,k}$, where ${\gamma^{\prime}}^{j}_{i,k}\subset B^{\prime}_{{n^{\prime}}^{j}_{i,k}}$ and ${\beta^{\prime}}_{i}^{j}$ is initiated with the rule $\delta^{j}_{i}(0):={n^{\prime}}^{j}_{i,0}-{n^{j}}_{i,0}=0$. We claim that $\beta_{i}^{j}$ meets the gluing curve if and only if ${\beta^{\prime}}_{i}^{j}$ intersects a gluing curve. Based on observations from [Bus86], we see that if $\delta^{j}_{i}(k)=0$ or $4$, crossing a side of length $c$ or $b$ does not change $\delta^{j}_{i}$ and crossing a side of length $a$ changes $\delta^{j}_{i}$ to $\delta^{j}_{i}+4$. Thus, $\delta^{j}_{i}(k)=0$ or $4$ for all $0\leq k\leq L$; that is, ${n^{j}}_{i,k}$ and ${n^{\prime}}^{j}_{i,k}$ differ by either 00 or $4$ (see, for example, any of the previous 3 figures). Then $\beta_{i}^{j}$ meets a gluing curve if and only if ${\beta^{\prime}}_{i}^{j}$ does. This establishes a natural bijection between $\{\beta^{j}_{i}\}_{j=1}^{2}$ and $\{{\beta^{\prime}}^{j}_{i}\}_{j=1}^{2}$, so one can make the same computations as before for calculating $C_{2}(\gamma)$, which counts the number of identical copies of $\gamma$ in $X_{2}$. We obtain in the end that $C_{1}(\gamma)=C_{2}(\gamma)$. One can therefore bijectively map copies of $\gamma$ in $X_{1}$ to those in $\gamma$ in $X_{2}$. ∎
@@ 3.2 Non-homeomorphic, iso-length-spectral surface amalgams
Recall that if two metric spaces are homeomorphic, then there is a quasi-isometry between their universal covers by the Milnor-Schwarz Theorem. From Proposition 3.4, it then follows that the chambers of two homeomorphic simple, thick surface amalgams are necessarily in bijective correspondence with each other. With this in mind, we now construct two isopectral, non-homeomorphic hyperbolic surface amalgams.
@@
Consider $S_{1}$ and $S_{2}$ from [Bus86] which have sets of systoles $\{\gamma_{i}\}_{i=1}^{4}$ and $\{\gamma^{\prime}_{i}\}_{i=1}^{4}$. Construct $X_{1}$ by identifying all the systoles $\gamma_{i}\subset X_{1}$. Similarly, glue together all the $\gamma^{\prime}_{i}\subset S_{2}$ to construct $X_{2}$.
@@
$X_{1}$ and $X_{2}$ are iso-length-spectral. We follow a similar strategy as before, by showing that the number of closed geodesics identical to a given closed geodesic $\gamma\subset X_{1}$ is the same in $X_{1}$ and $X_{2}$. Again, we decompose $\gamma$ into a union $\bigcup\limits_{i=1}^{N}\beta_{i}$ of geodesic segments that project to connected geodesic segments in $S_{1}$ beginning and ending on the gluing curve and count the number of ways to concatenate admissible translates of each $\beta_{i}$.
We will show that in $X_{1}$, there are four translates of $\beta_{i}$ with the same beginning and end points on the gluing curve. If $\beta_{i}$ begins on the right (resp. left) edge of a building block, we show there is one copy starting on the right edge (resp. left) of each even (resp. odd) labeled building block. In particular, we claim that each translated copy of $\beta_{i}$ will end on a building block with the same parity as the one $\beta_{i}$ ends on (and thus share an endpoint with $\beta_{i}$).
Let $\{\beta_{i}\}_{i=1}^{4}$ be the set of geodesic segments in $X_{1}$ which are translated copies of $\beta_{i}$ that either begin on the left or right sides of the odd or even labeled building blocks respectively, depending on where $\beta_{i}$ begins. Suppose $\beta_{i}^{1}:=\beta_{i}$. We check that $\beta_{i}^{j}$, where $j=2,3,$ or $4$, ends on a building block whose label has the same parity as the one $\beta_{i}^{1}$ ends on, allowing it to be concatenated with the next geodesic segment. For a fixed $j=J$, let $\beta_{i}^{J}=\bigcup\limits_{k=0}^{L}\gamma^{J}_{i,k}$, where each $\gamma^{J}_{k}$ is contained in a single building block, $B_{n^{J}_{k}}$. Define $d^{j}_{i}(k):=n_{i,k}^{j}-n_{i,k}^{1}$. Note that $d^{j}_{i}(0)$ is even since each $\beta_{i}^{j}$ must start on either an odd or even labeled building block. Crossing a side of length $c$ will not change $d^{j}_{i}(k)$. Furthermore, crossing a side of length $b$ (resp. $a$) will change (resp. not change) the parity of both $n_{i,k}^{j}$ and $n_{i,k}^{1}$, so the parity of $d^{j}_{i}(k)$ remains the same for all $0\leq k\leq L$. In particular, $d^{j}_{i}(L)$ is even, so the parity of $n_{i,L}^{j}$ will match that of $n_{i,1}^{j}$.
We can now compute $C_{1}(\gamma)$ using the same strategy as before. Using the same argument from Section 3.1 (replacing 1 and 2 with 3 and 4 respectively), we deduce that for each $12\leq i\leq N-$ 1:
$c_{i}$ :=#$:=\#\{\beta^{j}_{i}\text{ which can be concatenated with some fixed copy of }\beta_{i-1}\}$
=$=\begin{cases}3\text{ if $\beta_{i-1}$ ends on a right (resp. left) edge, and $\beta_{i}$ begins on a right (resp. left) edge;}\\ 4\text{ otherwise.}\end{cases}$
@@
By the same proof as before but replacing $0,1$ and $2$ with $2,3$, and $4$ respectively, we have $C_{1}(\gamma)=4\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(4)$ for Case 1, $C_{1}(\gamma)=4\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(3)$ for Case 2, and $C_{1}(\gamma)=\dfrac{4}{16}(4)\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(3)+\dfrac{12}{16}(4)\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(2)=\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(3)+3\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(2)$ for Case 3. The fractions from Case 3 come from the fact that there are now a total of 16 choices of pairs of admissible geodesic segments $\beta_{1}^{j}$ and $\beta_{N-1}^{k}$ where $j,k\in\{1,2,3,4\}$. For exactly four of these pairs, there is exactly one inadmissible copy of $\beta_{N}$ which has beginning point coinciding with the endpoint of $\beta_{N-1}^{k}$ and endpoint coinciding with the beginning point of $\beta_{1}^{j}$. For the 12 other pairs, there are two inadmissible copies of $\beta_{N}$.
We now show $C_{2}(\gamma)=C_{1}(\gamma)$. First, we claim there are four transplanted copies of $\beta_{i}$ that begin on the right (resp. left) side of each odd (resp. even) labeled building block and end at the same point. There is a natural isometry between gluing geodesics in $X_{1}$ and those in $X_{2}$; if a gluing geodesic $c_{n}$ is the right edge of $B_{n}$ and the left edge of $B_{n+1}$, let $\varphi(c_{n})$ be the right edge of $B_{n+1}$ and the left edge of $B_{n+2}$. If each $\beta_{i}^{j}$ begins at $x\in X_{1}$, we construct a set $\{{\beta^{\prime}}_{i}^{j}\}_{j=1}^{4}=\bigg{\{}\bigcup\limits_{k=0}^{L}{\gamma^{\prime}}^{j}_{i,k}\bigg{\}}_{j=1}^{4}$ of transplanted copies of $\beta_{i}^{j}$ which begin at $\varphi(x)\in X_{2}$ and pass through building blocks $B^{\prime}_{{n^{\prime}}_{i,k}^{j}}$. Initiate each ${\beta^{\prime}}_{i,k}^{j}$ with the rule $\delta^{j}_{i}(0):={n^{\prime}}_{i,0}^{j}-n_{i,0}^{j}=1$. By [Bus86], if $\delta^{j}_{i}=\pm 1$, crossing sides of length $a$ or $c$ do not affect $\delta^{j}_{i}$ while crossing a side of length $b$ sends ±$\pm 1$ to $\mp 1$. Regardless of the sequences of edges crossed, ${\delta^{\prime}}^{j}_{i}(L)=1$ or $-1$ for all $j$. Thus, each ${\beta^{\prime}}_{i}^{j}$ will end on the same side of a gluing geodesic that each $\beta_{i}^{j}$ ends on. Each ${\beta^{\prime}}_{i}^{j}$ will also end on a distinct building block.
@@
We now prove Theorem 1.3, our second main result. Consider, again, the surfaces from [Bus92]. We will glue three copies of each surface together to create $X_{1}$ and $X_{2}$. For both $X_{1}$ and $X_{2}$, there will be one gluing curve which will consist of unions of perpendiculars between edges of length $b$ that bisect all the building blocks except except ones labeled $6$ and $7$.
Note that in each surface, three pairs of perpendiculars form geodesics: the ones in building blocks with labels $1$ and $4$, $2$ and $3$, and $5$ and $8$. As a result, we need to specify how the geodesics are glued together. We will identify the halves of geodesics on the odd numbered building blocks and halves of geodesics on even numbered building blocks.
$X_{1}$ and $X_{2}$ are weak length isospectral. We use Buser’s transplantation technique. It suffices to consider closed geodesics that are not closed geodesics in copies of $S_{1}$ and $S_{2}$. We decompose $\gamma\subset X_{1}$ into a union of geodesic segments $\bigcup\limits_{j=1}^{N}\gamma_{j}$ each of which is contained entirely within a building block $B_{n_{j}}$. We must specify an algorithm to construct a transplanted copy of $\gamma$, $\gamma^{\prime}\subset X_{2}$, which is a union of geodesic segments $\bigcup\limits_{j=1}^{N}\gamma^{\prime}_{j}$ that project to connected components in $S_{2}$, start and end on the gluing curve in $X_{2}$, and do not backtrack. We will label the copies of $S_{1}$ in $X_{1}$ with $S_{1}^{1}$, $S_{1}^{2}$, and $S_{1}^{3}$ and similarly label the copies of $S_{2}$ in $X_{2}$. In the following algorithm, we will write $\delta^{s}(j)={n^{\prime}}^{s}_{j}-n^{s}_{j}$, which gives instructions for transplanting geodesic segments from $S_{1}^{s}$ to $S_{2}^{s}$.
@@
    If $\gamma_{j}$ lies in a building block indexed by an element in the set $\{1,4,5,8\}$, we set $\delta^{s}(j)=0$ if #$\#b$ is even and $4$ if #$\#b$ is odd.
@@
We now show that following (1), (2) and (3) from Algorithm 4.1, for each $\beta_{j,j+L}\subset S_{1}^{s}$, we can obtain a geodesic segment $\beta^{\prime}_{j,j+L}\in S_{2}^{s}$ which begins and ends on the gluing curve. If $\gamma_{j}\subset B_{n_{j}}$ for some $n_{j}\in\{1,4,5,8\}$, then we set $\delta^{s}(j)=0$ if #$\#b$ is even and $\delta^{s}(j)=4$ if #$\#b$ is odd so that regardless, $\gamma^{\prime}_{j}\subset B^{\prime}_{n^{\prime}_{j}}$, where $n^{\prime}_{j}=k+4\in\{1,4,5,8\}$. By [Bus86], if $\delta^{s}=0$ or $4$, crossing a side of length $a$ or $c$ leaves $\delta^{s}$ invariant, while crossing a side of length $b$ replaces $\delta^{s}$ by $\delta+4$. Thus, $\delta^{s}(j+L)=0$, so $B^{s}_{{n^{\prime}}_{(j+L)}}$ intersects a gluing curve since $B^{s}_{n_{(j+L)}}$ does. Suppose $\gamma_{j+L}\subset B_{k}$ where $k\in\{1,4,5,8\}$. Then setting $\delta^{s}(j)=0$ ensures that $\gamma^{\prime}_{j}$ begins on a building block intersecting a gluing curve and that $\delta^{s}(j+L)$ is 00 or $4$. Thus, $\gamma^{\prime}_{j+L}$ lies in a building block indexed by an element of the set $\{1,4,5,8\}$, which necessarily intersects the gluing curve.
This leaves the case ($\gamma_{j}\subset B_{n_{j}}$, where $n_{j}\in\{2,3\}$) and ($\gamma_{j+L}\subset B_{n_{j+L}}$, where $n_{j+L}\in\{2,3\}$). We set $\delta^{s}(j)=2$. Since $n_{j}\in\{2,3\}$, $n^{\prime}_{j}\in\{1,4,5,8\}$ and the building blocks indexed by this set all intersect the gluing curve. Buser observes that if $\delta=\pm 2$, crossing a side of length $a$ or $b$ replaces $\delta$ by $\delta+4$, while as before, crossing a side of length $c$ does not change $\delta$. Thus, if #$\#a+\#b$ is even (resp. odd), then $\delta^{s}(j+L)=2$ (resp. $-2$). Either way, since $n^{s}_{j+L}\in\{2,3\}$, we have that ${n^{\prime}}^{s}_{j+L}\in\{1,4,5,8\}$, and the building blocks indexed by this set all intersect the gluing curve.
@@
Consider a six-sheeted cover of $X_{1}$, $\widehat{X_{1}}$, which consists of six copies of a genus four surface with six boundary components (three-sheeted covers of $S_{2,2}$) and 12 copies of a torus with six boundary components (three-sheeted covers of $S_{1,2}$). Consider also six-sheeted cover of $X_{2}$, $\widehat{X_{2}}$, which consists of nine copies of genus three surfaces with six boundary components (double covers of $S_{2,3}$) and nine copies of a torus with six boundary components (six three-sheeted covers of $S_{1,2}$ and three six-sheeted covers of $S_{1,1}$) (see Figure 12). For both $\widehat{X_{1}}$ and $\widehat{X_{2}}$, there are six gluing curves, and exactly one boundary component in each chamber is glued to each gluing curve. Note that the Euler characteristics of the chambers in $\widehat{X_{1}}$, in ascending order, are $\{\underbrace{-12,...,-12}_{\times 6},\underbrace{-6,...,-6}_{\times 12}\}$ while those in $\widehat{X_{2}}$ are $\{\underbrace{-10,...,-10}_{\times 9},\underbrace{-6,...,-6}_{\times 9}\}$. Let $\{S_{i}\}$ and $\{T_{i}\}$ denote the collections of chambers of $\widehat{X_{1}}$ and $\widehat{X_{2}}$ respectively, labeled so that $\chi(S_{1})\leq\chi(S_{2})\leq...\leq\chi(S_{18})$ and $\chi(T_{1})\leq\chi(T_{2})\leq...\leq\chi(T_{18})$. A generalization of Proposition 3.3.2 of [Sta17], proved in Section 5.2 of [DST18], implies that $\pi_{1}(\widehat{X_{1}})$ and $\pi_{1}(\widehat{X_{2}})$ are abstractly commensurable (e.g. they do not have isomorphic finite-index subgroups) if and only if all the $\dfrac{\chi(S_{i})}{\chi(T_{i})}$ are equal. Note, however, that $\dfrac{\chi(S_{1})}{\chi(T_{1})}=\dfrac{-12}{-10}\neq\dfrac{-6}{-6}=\dfrac{\chi(S_{18})}{\chi(T_{18})}$. It then follows that $\pi_{1}(\widehat{X_{1}})$ and $\pi_{1}(\widehat{X_{2}})$ are not abstractly commensurable.
Since abstract commensurability is an equivalence relation, $\pi_{1}(X_{1})$ and $\pi_{1}(X_{2})$ also cannot be abstractly commensurable since $\pi_{1}(X_{i})$ is abstractly commensurable to $\pi_{1}(\widehat{X_{i}})$ for $i=1,2$. By the Galois correspondence of covering spaces for CW complexes, it then follows that $X_{1}$ and $X_{2}$ also cannot be commensurable, as claimed. ∎

### v2

@@ 3.1 Homeomorphic, non-isometric surface amalgams
Figure 7: An illustration for Case 3. Again, $N=3$ and $c_{2}=2$, so $C_{1}(\gamma)=2(2)=C_{2}(\gamma)$. We can check that there are indeed four identical, non-backtracking closed geodesics in each of $X_{1}$ and $X_{2}$.
We now construct an identical transplanted copy of each $\beta_{i}^{j}$, ${\beta^{\prime}}_{i}^{j}=\bigcup\limits_{k=0}^{L}{\gamma^{\prime}}^{j}_{i,k}$, where ${\gamma^{\prime}}^{j}_{i,k}\subset B^{\prime}_{{n^{\prime}}^{j}_{i,k}}$ and ${\beta^{\prime}}_{i}^{j}$ is initiated with the rule $\delta^{j}_{i}(0):={n^{\prime}}^{j}_{i,0}-{n}^{j}_{i,0}=0$ (see Equation 1). We claim that $\beta_{i}^{j}$ meets the gluing curve if and only if ${\beta^{\prime}}_{i}^{j}$ intersects a gluing curve. By Lemma 2.3, if $\delta^{j}_{i}(k)=0$ or $4$, crossing a side of length $c$ or $b$ does not change $\delta^{j}_{i}$ and crossing a side of length $a$ changes $\delta^{j}_{i}$ to $\delta^{j}_{i}+4$. Thus, $\delta^{j}_{i}(k)=0$ or $4$ for all $0\leq k\leq L$; that is, $n^{j}_{i,k}$ and ${n^{\prime}}^{j}_{i,k}$ differ by either 0 or $4$ (see, for example, any of the previous 3 figures). Then $\beta_{i}^{j}$ meets a gluing curve if and only if ${\beta^{\prime}}_{i}^{j}$ does. This establishes a natural bijection between $\{\beta^{j}_{i}\}_{j=1}^{2}$ and $\{{\beta^{\prime}}^{j}_{i}\}_{j=1}^{2}$, so one can make the same computations as before for calculating $C_{2}(\gamma)$, which counts the number of identical copies of $\gamma$ in $X_{2}$. We obtain in the end that $C_{1}(\gamma)=C_{2}(\gamma)$. One can therefore bijectively map copies of $\gamma$ in $X_{1}$ to those in $X_{2}$. ∎
@@ 3.2 Non-homeomorphic, iso-length-spectral surface amalgams
Recall that if two compact metric spaces are homeomorphic, then there is a quasi-isometry between their universal covers by the Milnor-Schwarz Lemma. From Proposition 3.4, it then follows that the chambers of two homeomorphic simple, thick surface amalgams are necessarily in bijective correspondence with each other. With this in mind, we now construct two isopectral, non-homeomorphic hyperbolic surface amalgams.
@@
Consider $S_{1}$ and $S_{2}$ from [Bus86] which have sets of systoles $\{\gamma_{i}\}_{i=1}^{4}$ and $\{\gamma^{\prime}_{i}\}_{i=1}^{4}$. Construct $X_{1}$ by identifying all the systoles $\gamma_{i}\subset X_{1}$ according to the orientations specified in Figure 9. Similarly, glue together all the $\gamma^{\prime}_{i}\subset S_{2}$ to construct $X_{2}$.
@@
$X_{1}$ and $X_{2}$ are iso-length-spectral. We follow a similar strategy as before, by showing that the number of closed geodesics identical to a given closed geodesic $\gamma\subset X_{1}$ is the same in $X_{1}$ and $X_{2}$. Again, we focus our attention on closed geodesics that do not intersect the gluing curve. We decompose $\gamma$ into a union $\bigcup\limits_{i=1}^{N}\beta_{i}$ of geodesic segments that project to connected geodesic segments in $S_{1}$ beginning and ending on the gluing curve and count the number of ways to concatenate admissible translates of each $\beta_{i}$.
We will show that in $X_{1}$, there are four translates of $\beta_{i}$ with the same beginning and end points on the gluing curve. If $\beta_{i}$ begins on the right (resp. left) edge of a building block, we show there is one copy starting on the right edge (resp. left) of each odd (resp. even) labeled building block. In particular, we claim that each translated copy of $\beta_{i}$ will end on a building block with the same parity as the one $\beta_{i}$ ends on (and thus share an endpoint with $\beta_{i}$).
Let $\{\beta^{j}_{i}\}_{j=1}^{4}$ be the set of geodesic segments in $X_{1}$ which are translated copies of $\beta_{i}$ that either all begin on the left sides of the even labeled building blocks or all begin on the right sides of the odd labeled building blocks, depending on where $\beta_{i}$ begins. Suppose $\beta_{i}^{1}:=\beta_{i}$. We check that $\beta_{i}^{j}$, where $j=2,3,$ or $4$, ends on a building block whose label has the same parity as the one $\beta_{i}^{1}$ ends on, allowing it to be concatenated with the next geodesic segment. Let $\beta_{i}^{j}=\bigcup\limits_{k=0}^{L}\gamma^{j}_{i,k}$, where each $\gamma^{j}_{k}$ is contained in a single building block, $B_{n^{j}_{k}}$. Define $d^{j}_{i}(k):=n_{i,k}^{j}-n_{i,k}^{1}$. Note that $d^{j}_{i}(0)$ is even since each $\beta_{i}^{j}$ must start on either an odd or even labeled building block. Crossing a side of length $c$ will not change $d^{j}_{i}(k)$. Furthermore, crossing a side of length $b$ (resp. $a$) will change (resp. not change) the parity of both $n_{i,k}^{j}$ and $n_{i,k}^{1}$, so the parity of $d^{j}_{i}(k)$ remains the same for all $0\leq k\leq L$. In particular, $d^{j}_{i}(L)$ is even, so the parity of $n_{i,L}^{j}$ will match that of $n_{i,1}^{j}$.
We can now compute $C_{1}(\gamma)$ using the same strategy as before. Using the same argument from Section 3.1 (replacing 1 and 2 with 3 and 4 respectively), we deduce that for each $2\leq i\leq N-1$:
[…]
@@
By the same proof as before but replacing $0,1$ and $2$ with $2,3$, and $4$ respectively, we have $C_{1}(\gamma)=4\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(4)$ for Case 1, $C_{1}(\gamma)=4\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(3)$ for Case 2, and $C_{1}(\gamma)=\dfrac{4}{16}(4)\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(3)+\dfrac{12}{16}(4)\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(2)=\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(3)+3\bigg{(}\prod\limits_{i=2}^{N-1}c_{i}\bigg{)}(2)$ for Case 3. The fractions from Case 3 come from the fact that there are now a total of 16 choices of pairs of geodesic segments $\beta_{1}^{j}$ and $\beta_{N-1}^{k}$ where $j,k\in\{1,2,3,4\}$. For exactly four of these pairs, there is exactly one inadmissible copy of $\beta_{N}$ which begins on the same building block as the end of $\beta_{N-1}^{k}$ and ends on the same building block as the beginning of $\beta_{1}^{j}$. For the 12 other pairs, there are two inadmissible copies of $\beta_{N}$.
We now show $C_{2}(\gamma)=C_{1}(\gamma)$. First, we claim there are four transplanted copies of $\beta_{i}$ that begin on the right (resp. left) side of each odd (resp. even) labeled building block and end at the same point. There is a natural isometry between gluing geodesics in $X_{1}$ and those in $X_{2}$; if a gluing geodesic $c_{n}$ is the right edge of $B_{n}$ and the left edge of $B_{n+1}$, let $\varphi(c_{n})$ be the right edge of $B_{n+1}$ and the left edge of $B_{n+2}$. If each $\beta_{i}^{j}$ begins at $x\in X_{1}$, we construct a set $\{{\beta^{\prime}}_{i}^{j}\}_{j=1}^{4}=\bigg{\{}\bigcup\limits_{k=0}^{L}{\gamma^{\prime}}^{j}_{i,k}\bigg{\}}_{j=1}^{4}$ of transplanted copies of $\beta_{i}^{j}$ which begin at $\varphi(x)\in X_{2}$ and pass through building blocks $B^{\prime}_{{n^{\prime}}_{i,k}^{j}}$. Initiate each ${\beta^{\prime}}_{i,k}^{j}$ with the rule $\delta^{j}_{i}(0):={n^{\prime}}_{i,0}^{j}-n_{i,0}^{j}=1$. By Lemma 2.3, if $\delta^{j}_{i}(k)=\pm 1$, crossing sides of length $a$ or $c$ does not affect $\delta^{j}_{i}(k)$ while crossing a side of length $b$ sends $\pm 1$ to $\mp 1$. Regardless of the sequences of edges crossed, ${\delta^{\prime}}^{j}_{i}(L)=1$ or 1-1 for all $j$. Thus, each ${\beta^{\prime}}_{i}^{j}$ will end on the same side of a gluing geodesic that each $\beta_{i}^{j}$ ends on. Each ${\beta^{\prime}}_{i}^{j}$ will also end on a distinct building block.
@@
We now prove Theorem 1.3, our second main result following the construction below.
Construction 4.1.
Consider, again, the surfaces from [Bus86]. We will glue three copies of each surface together to create $X_{1}$ and $X_{2}$. For both $X_{1}$ and $X_{2}$, there will be one gluing curve which will consist of unions of perpendiculars between edges of length $b$ that bisect all the building blocks except ones labeled $6$ and $7$.
Note that in each surface, three pairs of perpendiculars form geodesics: the ones in building blocks with labels $1$ and $4$, 22 and $3$, and $5$ and $8$. As a result, we need to specify how the geodesics are glued together. We will identify the halves of geodesics on the odd numbered building blocks and halves of geodesics on even numbered building blocks. See Figure 10 for the orientations of the gluing curves.
$X_{1}$ and $X_{2}$ are weak length isospectral. We use Buser’s transplantation technique. As before, it suffices to consider closed geodesics that are not closed geodesics in copies of $S_{1}$ and $S_{2}$. We decompose $\gamma\subset X_{1}$ into a union of geodesic segments $\bigcup\limits_{j=1}^{N}\gamma_{j}$ each of which is contained entirely within a building block $B_{n_{j}}$. We must specify an algorithm to construct a transplanted copy of $\gamma$, $\gamma^{\prime}\subset X_{2}$, which is a union $\bigcup\limits_{k=1}^{N}\gamma^{\prime}_{k}$ of geodesic segments that are each completely contained within a building block in $S_{2}$, start and end on the gluing curve in $X_{2}$, and do not backtrack. We will label the copies of $S_{1}$ in $X_{1}$ with $S_{1}^{1}$, $S_{1}^{2}$, and $S_{1}^{3}$ and similarly label the copies of $S_{2}$ in $X_{2}$. In the following algorithm, we will write $\delta^{s}(j)={n^{\prime}}^{s}_{j}-n^{s}_{j}$, which gives instructions for transplanting geodesic segments from $S_{1}^{s}$ to $S_{2}^{s}$.
@@
    If $\gamma_{j}$ lies in a building block indexed by an element in the set $\{1,4,5,8\}$, we set $\delta^{s}(j)=0$ if #b\#b is even and $4$ if #b\#b is odd.
@@
We now show that following (1), (2) and (3) from Algorithm 4.2, for each $\beta_{j,j+L}\subset S_{1}^{s}$, we can obtain a geodesic segment $\beta^{\prime}_{j,j+L}\in S_{2}^{s}$ which begins and ends on the gluing curve. If $\gamma_{j}\subset B_{n_{j}}$ for some $n_{j}\in\{1,4,5,8\}$, then we set $\delta^{s}(j)=0$ if #b\#b is even and $\delta^{s}(j)=4$ if #b\#b is odd so that regardless, $\gamma^{\prime}_{j}\subset B^{\prime}_{n^{\prime}_{j}}$, where $n^{\prime}_{j}=k+4\in\{1,4,5,8\}$. By Lemma 2.3, if $\delta^{s}=0$ or $4$, crossing a side of length $a$ or $c$ leaves $\delta^{s}$ invariant, while crossing a side of length $b$ replaces $\delta^{s}$ by $\delta+4$. Thus, $\delta^{s}(j+L)=0$, so $B^{s}_{{n^{\prime}}_{(j+L)}}$ intersects a gluing curve since $B^{s}_{n_{(j+L)}}$ does. Suppose $\gamma_{j+L}\subset B_{k}$ where $k\in\{1,4,5,8\}$. Then setting $\delta^{s}(j)=0$ ensures that $\gamma^{\prime}_{j}$ begins on a building block intersecting a gluing curve and that $\delta^{s}(j+L)$ is 0 or $4$. Thus, $\gamma^{\prime}_{j+L}$ lies in a building block indexed by an element of the set $\{1,4,5,8\}$, which necessarily intersects the gluing curve.
This leaves the case ($\gamma_{j}\subset B_{n_{j}}$, where $n_{j}\in\{2,3\}$) and ($\gamma_{j+L}\subset B_{n_{j+L}}$, where $n_{j+L}\in\{2,3\}$). We set $\delta^{s}(j)=2$. Since $n_{j}\in\{2,3\}$, $n^{\prime}_{j}\in\{1,4,5,8\}$ and the building blocks indexed by this set all intersect the gluing curve. By Lemma 2.3, $\delta=\pm 2$, crossing a side of length $a$ or $b$ replaces $\delta$ by $\delta+4$, while as before, crossing a side of length $c$ does not change $\delta$. Thus, if #a+#b\#a+\#b is even (resp. odd), then $\delta^{s}(j+L)=2$ (resp. 2-2). Either way, since $n^{s}_{j+L}\in\{2,3\}$, we have that ${n^{\prime}}^{s}_{j+L}\in\{1,4,5,8\}$, and the building blocks indexed by this set all intersect the gluing curve.
@@
[…]